A Christmas tree survey was conducted to determine there was a difference between urban rural tree users in terms of artificial tree use. The results are presented in the table below. What is the 90% CI for the difference between the proportions (urban minus rural) of artificial tree users?

Total Tree Users Number with Artificial Trees 89 64 261 160 Urban RuralP - ל± 1.645 n2

a.

(-0.211, 0.129)

b.

(0.023, 0.441)

c.

(-0.139, 0.021)

d.

(-0.246, 0.234)
Answer #1

Sample proportion \hat{p}1 = 89 / 261 = 0.341

Sample proportion \hat{p}2 = 64 / 160 = 0.400

90% confidence interval for p1 - p2 is

(\hat{p}1 - \hat{p}2) - 1.945 * sqrt[ \hat{p}1 ( 1 - \hat{p}1) / n1 + \hat{p}2 ( 1 - \hat{p}2) / n2 ) ] < p1 - p2 <

(\hat{p}1 - \hat{p}2) + 1.945 * sqrt[ \hat{p}1 ( 1 - \hat{p}1) / n1 + \hat{p}2 ( 1 - \hat{p}2) / n2 ) ]

(0.341 - 0.4) - 1.645 * sqrt [ 0.341 * 0.659 / 261 + 0.4 * 0.6 / 160 ] < p1 - p2 <

(0.341 - 0.4) + 1.645 * sqrt [ 0.341 * 0.659 / 261 + 0.4 * 0.6 / 160 ]

-0.139 < p1 - p2 < 0.021

90% CI is ( -0.139 , 0.021 )